OpenStudy (anonymous):

How do I get all the solutions for x^8-256=0

OpenStudy (perl):

you want 8 solutions?

OpenStudy (perl):

that equation has 2 real solutions , x ={2, -2}. and 6 non real solutions

OpenStudy (anonymous):

You get \[ x^8-256=(x-2) (x+2) \left(x^2+4\right) \left(x^4+16\right) \] You can finish it from here to find the 6 complex roots

OpenStudy (anonymous):

In addition to the real ones 2 and -2